OpenStudy (anonymous):

Which expressions could complete this equation so that it has one solution? 4(4x-3)-6x Choose exactly two answers that are correct. A. 12x-12-02x B. 9x-12 C. 4x+2(3x-6) D. 4(2x+3)-24

OpenStudy (anonymous):

I'm trying to get different help. Thank you though.

pooja195 (pooja195):

I think that help was decent

OpenStudy (anonymous):

It was confusing.

pooja195 (pooja195):

You are looking for answers.This site is meant for HELP not answers sorry.

OpenStudy (anonymous):

@pooja195 your the best help i've had.

pooja195 (pooja195):

If you are willing to learn i'll help :) 4(4x-3)-6x Start by distribtuing 4(4)=? 4(-3)=?

OpenStudy (anonymous):

4(4)=16 4(-3)=-12

OpenStudy (anonymous):

Right?

pooja195 (pooja195):

Right so we have 16x-12-6x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why did you delete that comment?

pooja195 (pooja195):

4x+2(3x-6) 4x+6x=12 10x-12 <---match your options with this

pooja195 (pooja195):

From the Question 16x-12-6x 10x-12 C is one of the choices because it is equivalent to this line "10x-12" 4x+2(3x-6) 4x+6x=12 10x-12 Find the other answer

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Or a

pooja195 (pooja195):

Tell me why you think those are possible answers

OpenStudy (anonymous):

Because D goes along the lines of what we are doing. Idk why I said A

pooja195 (pooja195):

4(2x+3)-24 8x+12-24 Is this equivalent to 10x-12 ?

pooja195 (pooja195):

?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

its not.

OpenStudy (anonymous):

i know.

OpenStudy (anonymous):

Alright. serious now

pooja195 (pooja195):

ok we know D is not the answer. Try again.

OpenStudy (anonymous):

It's b

pooja195 (pooja195):

Nope 9x-12 is not equal to 10x-12

OpenStudy (anonymous):

its a and c

pooja195 (pooja195):

Well done =)